# SMMILE: An Expert-Driven Benchmark for Multimodal Medical In-Context Learning

### v1

@@ Abstract
Multimodal in-context learning (ICL) remains underexplored despite significant potential for domains such as medicine. Clinicians routinely encounter diverse, specialized tasks requiring adaptation from limited examples, such as drawing insights from a few relevant prior cases or considering a constrained set of differential diagnoses. While multimodal large language models (MLLMs) have shown advances in medical visual question answering (VQA), their ability to learn multimodal tasks from context is largely unknown. We introduce SMMILE, the first expert-driven multimodal ICL benchmark for medical tasks. Eleven medical experts curated problems, each including a multimodal query and multimodal in-context examples as task demonstrations. SMMILE encompasses 111 problems (517 question-image-answer triplets) covering 6 medical specialties and 13 imaging modalities. We further introduce SMMILE++, an augmented variant with 1038 permuted problems. A comprehensive evaluation of 15 MLLMs demonstrates that most models exhibit moderate to poor multimodal ICL ability in medical tasks. In open-ended evaluations, ICL contributes only 8% average improvement over zero-shot on SMMILE and 9.4% on SMMILE++. We observe a susceptibility for irrelevant in-context examples: even a single noisy or irrelevant example can degrade performance by up to 9.5%. Moreover, example ordering exhibits a recency bias, i.e., placing the most relevant example last can lead to substantial performance improvements by up to 71%. Our findings highlight critical limitations and biases in current MLLMs when learning multimodal medical tasks from context. SMMILE is available at https://smmile-benchmark.github.io.
@@ 1 Introduction
In-context learning (ICL) has been widely studied as the striking ability of large language models (LLMs) to learn and generalize to new tasks at inference time when provided with a few demonstration examples in their input context, without requiring any parameter updates (brown2020language, ). Given a set of relevant labeled examples in the input prompt, ICL enables models to flexibly adapt to provided contexts, contributing to applications like retrieval-augmented generation (karpukhin2020dense, ; lewis2020retrieval, ; zakka2024almanac, ; xiong2024benchmarking, ) and model personalization.
@@
Although ICL has been predominantly studied in the context of LLMs, recent works have explored extensions of ICL to multimodal settings (alayrac2022flamingo, ; laurenccon2023obelics, ; laurenccon2024building, ; laurenccon2024matters, ). Multimodal ICL holds particular promise for the domain of medicine due to the close parallels between ICL and clinical workflows; in real-world medical settings, clinicians are routinely asked to address specialized tasks given knowledge of limited prior examples, such as a few relevant prior cases or a constrained set of differential diagnoses. Models capable of performing multimodal ICL in high-stakes medical settings must be carefully assessed for reliability. Although some prior works have proposed strategies for evaluating the ICL capabilities of multimodal LLMs (MLLMs) in the general domain (zeng2024mllmsperformtexttoimageincontext, ; Baldassini_2024_CVPR, ; chen2024multimodallargelanguagemodels, ; zong2024vl, ), no benchmarks have been previously developed to systematically evaluate multimodal ICL in the medical domain. Additionally, existing few-shot evaluations in medical settings often randomly select examples rather than focusing on specific task demonstrations (moor2023med, ; yan-etal-2023-style, ), which may partially explain why minimal improvements over zero-shot evaluations are often observed.
In this work, we aim to address these challenges by introducing the Stanford Multimodal Medical In-context Learning (SMMILE) benchmark. Notably, SMMILE is an expert-driven benchmark, developed in collaboration with an international team of 11 medical experts. Our contributions are summarized below:
@@
  • We evaluate 15 MLLMs on our benchmarks, including both open-source and closed-source models with diverse architectures and model sizes. Our results show that existing MLLMs struggle to effectively learn from multimodal in-context examples in the medical setting, with ICL contributing to minimal performance boosts over zero-shot evaluations across most evaluated models. In open-ended settings, even the best-performing models (GPT-4o and Qwen2.5-VL-72B) are only capable of answering approximately half of the questions accurately. These results expose a significant shortcoming of current MLLMs: although in-context examples are manually designed to serve as effective task demonstrations, MLLMs are unable to accurately learn the task at hand.
  • The manually-curated and high-quality nature of the SMMILE benchmark can help reveal insights into how effective in-context examples can be selected for MLLMs. To this end, we perform an in-depth analysis of two critial factors associated with selecting in-context examples. First, we demonstrate that the quality of in-context examples is important: the inclusion of just one irrelevant sample in the in-context example list can impair performance. Second, we demonstrate that the order of in-context examples matters: all evaluated MLLMs suffer from recency bias, where placing the most relevant in-context examples later in the example list can lead to improved performance.
@@ 2.1 Dataset Curation
In order to collect data, we first recruited clinical experts to contribute multimodal ICL problems. In this setting, each problem consists of (1) a query to be posed to a MLLM, including a question, an associated non-text media item (e.g. an image), and a ground-truth answer; and (2) two or more in-context examples, each including a question, an associated non-text media item, and an answer. Examples are designed to serve as relevant task demonstrations that support a model in learning the task at hand. We provide sample queries and examples from SMMILE in Figure 1.
@@
The problem creation pipeline for SMMILE involves a guided, step-by-step workflow. First, the clinical expert is presented with a set of detailed instructions, which cover topic scope, data sourcing, and answer formatting (Appendix A.1). Next, the expert is directed to the homepage interface, where they initialize a new problem (Appendix A.2). Then, the problem creation tool is loaded (Appendix A.3), which enables the expert to select the relevant medical specialty as well as add, remove, or reorder panels for in-context examples and the final query. Finally, upon clicking “Submit,” the expert is shown an overview of the completed problem for validation or further editing (Appendix A.4). Our pipeline is designed to ensure consistency, adherence to guidelines, and ease of use through the problem creation process.
We then performed manual quality control, which involved the following three steps. First, each problem was manually inspected by two different authors to check for errors, irregularities, or other inaccuracies. Second, each problem was annotated and categorized as shown in Figure 3 (A-F). Third, all text was put through spell check software and manually adjusted when needed. This resulted in 15 grammar and spelling changes to questions and 6 grammar and spelling changes to answers. Additionally, 8 problems had to be modified to make exact match (EM) evaluations possible, including 1 phrasing change, 5 insertions of additional in-context examples, and 2 modifications to queries.
@@ 2.2 Benchmark Details
Figure 3 analyzes the composition of SMMILE with several descriptive statistics. As shown in Graphs A-D, the dataset is primarily composed of diagnostic and classification problems, with about three-quarters requiring free response formats. While many cases are common in clinical practice, over one-third represent uncommon presentations. The majority are rated as difficult for current LLMs. Graphs E and F demonstrate a diversity of image types, covering 6 medical specialties and 13 imaging modalities. Graph G summarizes the distribution of in-context examples per problem, and Graph H details the distribution of question and answer lengths across the dataset.
We leverage the SMMILE dataset to design two benchmark tasks: (1) open-ended generation, where a MLLM is presented with a query question and image and tasked with generating a free-text response, and (2) closed-ended generation, where the MLLM selects an answer from a closed set of possible choices obtained from the in-context example set. Additionally, we introduce a large-scale, augmented dataset called SMMILE++ by permuting in-context examples from a subset of problems in the original dataset. To find permutable problems, we excluded all reasoning problems. An upper limit of $4!=24$ permutations per problem was used, implying that problems with more than 4 in-context examples were shuffled until 24 unique permutations had been created. The final SMMILE++ dataset includes 1038 problems. Descriptive statistics for SMMILE++ are presented in Figure 9.
Figure 3: Dataset characteristics. (A–D) Distribution of four key categorical annotations across the unique problems: (A) answer format, (B) rarity of the clinical case based on how often clinicians would experience the medical concepts included in each problem, (C) primary cognitive process required (where reasoning classification is defined by final problem not having direct support in its in-context example set), and (D) rated difficulty for state-of-the-art (SOTA) LLMs based on model performance. Hard difficulty is assigned when tested SOTA LLMS are unable to answer the problem, otherwise medium difficulty is assigned to the problem. (E–F) Horizontal barplots showing the breakdown of each problem by its main medical specialty (E) and by main image type used (F). (G) Histogram of the number of in-context examples provided per problem. (H) Overlaid histograms of the character-length distributions for questions versus answers. All panels are based on the 111 problems included in SMMILE.
@@ 3.1 Experimental setup
We evaluate a total of 15 state-of-the-art MLLMs encompassing a range of model sizes (0.5B to 90B parameters for the open-source models), pretraining domains (general-purpose MLLMs and domain-specific medical MLLMs), access types (open-source and closed-source), and model architectures. Open-source models considered in this work include: LLaVA-v1.5 (7B and 13B) (liu2024improvedbaselinesvisualinstruction, ), LLaVA-NeXT-7B (liu2024llavanext, ), LLaVA-OneVision (0.5B and 7B) (li2024llavaonevisioneasyvisualtask, ), LLaVA-Med-7B (li2023llava, ), Llama-3.2-Vision-90B (grattafiori2024llama3herdmodels, ), MedVLM-R1 (pan2025medvlmr1incentivizingmedicalreasoning, ), MedGemma 4B (medgemma, ), and Qwen2.5-VL (3B, 7B, 32B, and 72B) (bai2025qwen25vltechnicalreport, ). In particular, LLaVA-Med-7B, MedGemma 4B, and MedVLM-R1 are domain-specific MLLMs designed specifically for medical tasks. Closed-source models considered in this work include GPT-4o (openai2024gpt4ocard, ) and Claude 3.7 Sonnet (claude, ). For all models, we use a standard input prompt consisting of a system message, in-context examples, and the query image and question. To ensure fair comparison, we set the maximum generation length to 512 tokens for all models across all open-ended tasks. All experiments with local models were conducted on a single-node compute server equipped with 8 NVIDIA H200 (141 GB) GPUs. For more details on computational resources, please refer to Appendix Section C.
@@
For open-ended evaluations, we evaluate MLLM-generated outputs using two primary metrics. First, the Exact Match (EM) metric counts a model generation as correct (score of 100) if it exactly matches the ground-truth answer, and incorrect (score of 0) otherwise. During evaluation, answers are normalized to account for minor variations in formatting, punctuation, and capitalization before comparison. Second, the LLM-as-a-Judge approach provides a text-only LLM (Llama3.3 70B) with both the model generation and the ground-truth answer; the model is then prompted to evaluate accuracy. This metric captures semantic correctness beyond exact string matching, making it particularly valuable for medical reasoning tasks where multiple phrasings might convey the same diagnosis or finding. The LLM provides a binary judgment (0 for incorrect, 1 for correct) for each generated output, and the final score represents the percentage of outputs judged as correct. For closed-ended (multiple-choice) evaluations, we simply measure the accuracy of selecting the correct option.
To estimate variability in our metrics, we employ a bootstrap resampling approach with $N_{\text{bootstrap}}=1000$ bootstrap iterations. For each iteration, we randomly sample with replacement from the original results to create a simulated dataset of the same size as the original dataset, then calculate the accuracy for this bootstrap sample. We report the mean accuracy and standard deviation across all 1000 bootstrap samples. We use a fixed random seed for reproducibility.
@@ 3.2 Benchmarking MLLMs with SMMILE
In Table 1, we report performance metrics resulting from evaluations of 15 state-of-the-art MLLMs on SMMILE. Several trends are observable from these results.
  • ICL shows mixed results with concerning baseline failures despite average improvements. While the average performance improvement across the 15 models in the open-ended LLM-as-a-Judge setting is 8.01% (absolute) and 31.2% (relative), this masks a troubling heterogeneity in ICL effectiveness. Notably, 7 out of 15 models perform worse than even a Random baseline (randomly selecting an answer from in-context examples, 27.86%): MedVLM-R1 (26.74%), LLaVA-OneVision-7B (24.25%), LLaVA-NeXT-7B (23.66%), LLaVA-OneVision-0.5B (21.63%), LLaVA-v1.5-13B (20.91%), LLaVA-v1.5-7B (18.727%), and LLaVA-Med-7B (10.19%). Some models even show performance degradation with ICL, such as LLaVA-Med-7B dropping by more than half from 21.65% to 10.19%. The average improvement is driven primarily by a few models showing substantial gains: LLaVA-NeXT-7B with 107.2% (11.42$\rightarrow$23.66%), Qwen2.5-VL-32B with 65.4% (25.27$\rightarrow$41.79%), and Qwen2.5-VL-72B with 42.4% (29.90$\rightarrow$42.59%) relative improvement, respectively. This highly variable performance reveals that ICL benefits remain model-specific and unreliable.
  • GPT-4o is the overall leader. With ICL, GPT-4o delivers the best open-ended score (49.88%) and the best closed-ended MCQA accuracy (58.85%), demonstrating the most effective multimodal reasoning capabilities across task formats.
@@
  • Closed-ended questions are easier for MLLMs. Five models achieve an accuracy greater than 50% on MCQA (GPT-4o, Claude 3.7 Sonnet, Llama-3.2-Vision-90B, Qwen2.5-VL-32B, Qwen2.5-VL-72B). A text-only Llama 3.3 baseline still achieves an accuracy of 38.6%, suggesting that multiple-choice items rely less on precise visual grounding than open-ended generation.
  • Model scale is not the sole determinant of success. Smaller Qwen (3-7B) and LLaVA (0.5-7B) variants lag well behind larger models. However, the Qwen2.5-VL-32B model outperforms its 72B counterpart on EM and approximately matches it on MCQA.
Table 1: We benchmark 15 MLLMs on SMMILE, reporting both zero-shot performance as well as performance with in-context examples. The best result is bolded for each task and evaluation metric. Text-only baseline used Llama 3.3 70B. ∗∗Llava-Med-7B refers to LLaVA-Med-v1.5-Mistral-7B.

### v2

@@ Abstract
Multimodal in-context learning (ICL) remains underexplored despite significant potential for domains such as medicine. Clinicians routinely encounter diverse, specialized tasks requiring adaptation from limited examples, such as drawing insights from a few relevant prior cases or considering a constrained set of differential diagnoses. While multimodal large language models (MLLMs) have shown advances in medical visual question answering (VQA), their ability to learn multimodal tasks from context is largely unknown. We introduce SMMILE, the first expert-driven multimodal ICL benchmark for medical tasks. Eleven medical experts curated problems, each including a multimodal query and multimodal in-context examples as task demonstrations. SMMILE encompasses 111 problems (517 question-image-answer triplets) covering 6 medical specialties and 13 imaging modalities. We further introduce SMMILE++, an augmented variant with 1038 permuted problems. A comprehensive evaluation of 15 MLLMs demonstrates that most models exhibit moderate to poor multimodal ICL ability in medical tasks. In open-ended evaluations, ICL contributes only an 8% average improvement over zero-shot on SMMILE and 9.4% on SMMILE++. We observe a susceptibility for irrelevant in-context examples: even a single noisy or irrelevant example can degrade performance by up to 9.5%. Moreover, we observe that MLLMs are affected by a recency bias, where placing the most relevant example last can lead to substantial performance improvements of up to 71%. Our findings highlight critical limitations and biases in current MLLMs when learning multimodal medical tasks from context. SMMILE is available at https://smmile-benchmark.github.io.
@@ 1 Introduction
In-context learning (ICL) has been widely studied as the striking ability of large language models (LLMs) to generalize to new tasks at inference time when provided with a few demonstration examples in their input context, without requiring any parameter updates (brown2020language, ). Given a set of relevant labeled examples in the input prompt, ICL enables models to flexibly adapt to provided contexts, contributing to applications like retrieval-augmented generation (karpukhin2020dense, ; lewis2020retrieval, ; zakka2024almanac, ; xiong2024benchmarking, ) and model personalization.
@@
Although ICL has been predominantly studied in the context of LLMs, recent works have explored extensions of ICL to multimodal settings (alayrac2022flamingo, ; laurenccon2023obelics, ; laurenccon2024building, ; laurenccon2024matters, ). Multimodal ICL holds particular promise for the domain of medicine due to the close parallels between ICL and clinical workflows; in real-world medical settings, clinicians are routinely asked to address specialized tasks given knowledge of limited prior examples, such as a few relevant prior cases or a constrained set of differential diagnoses. Models capable of performing multimodal ICL in high-stakes medical settings must be carefully assessed for reliability. Although some prior works have proposed strategies for evaluating the ICL capabilities of multimodal LLMs (MLLMs) in the general domain (zeng2024mllmsperformtexttoimageincontext, ; Baldassini_2024_CVPR, ; chen2024multimodallargelanguagemodels, ; zong2024vlbench, ), no benchmarks have been previously developed to systematically evaluate multimodal ICL in the medical domain. Additionally, existing few-shot evaluations in medical settings often randomly select examples rather than focusing on specific task demonstrations (moor2023med, ; yan-etal-2023-style, ), which may partially explain why minimal improvements over zero-shot evaluations are often observed.
In this work, we aim to address these challenges by introducing the Stanford Multimodal Medical In-context Learning (SMMILE) benchmark. Notably, SMMILE is an expert-driven benchmark, developed in collaboration with an international team of 11 medical experts. Our contributions are:
@@
  • We evaluate 15 MLLMs on our benchmarks, including both open-source and closed-source models with diverse architectures and model sizes. Model performance is assessed using both automated metrics as well as human expert evaluations. Our results show that existing MLLMs struggle to effectively learn from multimodal in-context examples in the medical setting, with ICL contributing to minimal performance boosts over zero-shot evaluations across most evaluated models. In open-ended settings, even the best-performing models (GPT-4o and Qwen2.5-VL-72B) are only capable of answering approximately half of the questions accurately. These results expose a significant shortcoming of current MLLMs: although in-context examples are manually designed to serve as effective task demonstrations, MLLMs are unable to accurately learn the task at hand.
  • The manually-curated and high-quality nature of the SMMILE benchmark can help reveal insights into how effective in-context examples can be selected for MLLMs. To this end, we perform an in-depth analysis of two critial factors associated with selecting in-context examples. First, we demonstrate that the quality of in-context examples is important: the inclusion of just one irrelevant sample in the in-context example list can impair performance. Second, we demonstrate that the order of in-context examples matters: all evaluated MLLMs suffer from recency bias, where placing the most relevant in-context examples later in the example list can improve performance.
@@
Related Work
Our work is motivated by prior research on in-context learning, medical MLLMs, and benchmarking efforts. We discuss related works in Appendix Section A.
@@ 2.1 Dataset Curation
In order to collect data, we first recruited clinical experts to contribute multimodal ICL problems. In this setting, each problem consists of (1) a query to be posed to a MLLM, including a question, an associated non-text media item (e.g. an image), and a ground-truth answer; and (2) two or more in-context examples, each including a question, an associated non-text media item, and an answer. Examples are designed to serve as relevant task demonstrations that support a model in learning the task at hand. We provide sample problems from SMMILE in Figure 1.
@@
The problem creation pipeline for SMMILE involves a guided, step‐by‐step workflow. First, the clinical expert is presented with a set of detailed instructions, which cover topic scope, data sourcing, and answer formatting (Appendix B.1). Next, the expert is directed to the homepage interface, where they initialize a new problem (Appendix B.2). Then, the problem creation tool is loaded (Appendix B.3), which enables the expert to select the relevant medical specialty as well as add, remove, or reorder panels for in‐context examples and the final query. Finally, upon clicking “Submit,” the expert is shown an overview of the completed problem for validation or further editing (Appendix B.4). Our pipeline is designed to ensure consistency, adherence to guidelines, and ease of use through the problem creation process.
We then performed manual quality control, which involved the following three steps. First, each problem was manually inspected by two different authors to check for errors, irregularities, or other inaccuracies. Second, each problem was annotated and categorized as shown in Figure 3 (A-F). Third, all text was put through spell check software and manually adjusted when needed. This resulted in 15 grammar and spelling changes to questions and 6 grammar and spelling changes to answers. Additionally, 8 problems had to be modified to make exact match (EM) evaluations possible, including 1 phrasing change, 5 insertions of additional in-context examples, and 2 query edits.
@@ 2.2 Benchmark Details
Figure 3 analyzes the composition of SMMILE with several descriptive statistics. As shown in Graphs A-C, the dataset is primarily composed of diagnostic and classification problems, with about three-quarters requiring free response formats. While many cases are common in clinical practice, over one-third represent uncommon presentations. Graph D characterizes problem difficulty, as rated by clinical experts during data curation. Specifically, experts considered: (1) whether the problem required complex multimodal reasoning beyond simple visual pattern matching, (2) the degree of specialized medical knowledge needed, and (3) to what extent successful answers would likely require leveraging the provided in-context examples rather than relying solely on pre-trained knowledge. Graphs E and F demonstrate a diversity of image types, covering 6 medical specialties and 13 imaging modalities. Graph G summarizes the distribution of in-context examples per problem, and Graph H details the distribution of question and answer lengths across the dataset.
We leverage the SMMILE dataset to design two benchmark tasks: (1) open-ended generation, where a MLLM is presented with a query question and image and tasked with generating a free-text response, and (2) closed-ended generation, where the MLLM selects an answer from a closed set of possible choices obtained from the in-context example set. Additionally, we introduce a large-scale, augmented dataset called SMMILE++ by permuting in-context examples from a subset of problems in the original dataset. To find permutable problems, we excluded all reasoning problems. An upper limit of $4!=24$ permutations per problem was used, implying that problems with more than 4 in-context examples were shuffled until 24 unique permutations had been created. The final SMMILE++ dataset includes 1038 problems. Descriptive statistics for SMMILE++ are presented in Appendix Section C.
Figure 3: Dataset characteristics. (A–D) Distribution of four key categorical annotations across the unique problems: (A) answer format, (B) rarity of the clinical case based on how often clinicians would experience the medical concepts included in each problem, (C) primary cognitive process required (where reasoning classification is defined by final problem not having direct support in its in-context example set), and (D) rated difficulty. (E–F) Horizontal barplots showing the breakdown of each problem by its main medical specialty (E) and by main image type used (F). (G) Histogram of the number of in-context examples provided per problem. (H) Overlaid histograms of the character-length distributions for questions versus answers. All panels are based on the 111 problems included in SMMILE.
@@ 3.1 Experimental setup
We evaluate a total of 15 state-of-the-art MLLMs encompassing a range of model sizes (0.5B to 90B parameters for the open-source models), pretraining domains (general-purpose MLLMs and domain-specific medical MLLMs), access types (open-source and closed-source), and model architectures. Open-source models considered in this work include: LLaVA-v1.5 (7B and 13B) (liu2024improvedbaselinesvisualinstruction, ), LLaVA-NeXT-7B (liu2024llavanext, ), LLaVA-OneVision (0.5B and 7B) (li2024llavaonevisioneasyvisualtask, ), LLaVA-Med-7B (li2023llava, ), Llama-3.2-Vision-90B (grattafiori2024llama3herdmodels, ), MedVLM-R1 (pan2025medvlmr1incentivizingmedicalreasoning, ), MedGemma 4B (medgemma, ), and Qwen2.5-VL (3B, 7B, 32B, and 72B) (bai2025qwen25vltechnicalreport, ). In particular, LLaVA-Med-7B, MedGemma 4B, and MedVLM-R1 are domain-specific MLLMs designed specifically for medical tasks. Closed-source models considered in this work include GPT-4o (openai2024gpt4ocard, ) and Claude 3.7 Sonnet (claude, ). For all models, we use a standard input prompt consisting of a system message, in-context examples, and the query image and question. To ensure fair comparison, we set the maximum generation length to 512 tokens for all models across all open-ended tasks.
@@
For open-ended evaluations, we evaluate MLLM-generated outputs using two metrics. First, the Exact Match (EM) metric counts a model generation as correct (score of 100) if it exactly matches the ground-truth answer, and incorrect (score of 0) otherwise. During evaluation, answers are normalized to account for minor variations in formatting, punctuation, and capitalization before comparison. Second, the LLM-as-a-Judge approach provides a text-only LLM (Llama3.3 70B) with both the model generation and the ground-truth answer; the model is then prompted to evaluate accuracy. The LLM provides a binary judgment (0 for incorrect, 1 for correct) for each generated output, and the final score represents the percentage of outputs judged as correct. For closed-ended (multiple-choice) evaluations, we measure the accuracy of selecting the correct option.
To estimate sampling variability in our metrics, we employ a bootstrap resampling approach with $N_{\text{bootstrap}}=1000$ bootstrap iterations. For each iteration, we randomly sample with replacement from the original results to create a simulated dataset of the same size as the original dataset, then calculate the accuracy for this bootstrap sample. We report the mean accuracy and standard deviation across all 1000 bootstrap samples. We use a fixed random seed for reproducibility.
To complement automated metrics, we conducted Human Expert evaluations, where five clinical experts independently evaluated model responses in both zero-shot and ICL settings. Each response was assessed by two different clinicians using binary ratings (correct/incorrect). Inter-rater agreement was perfect in the ICL setting (100%) and ranged from 98.2% to 100% in the zero-shot setting, demonstrating high reliability.
Additional experimental details are provided in Appendix Section D.
@@ 3.2 Benchmarking MLLMs with SMMILE
In Table 1, we report performance metrics resulting from evaluations of 15 state-of-the-art MLLMs on SMMILE. Several trends are observable from these results:
[…]
  • GPT-4o is the overall leader. With ICL, GPT-4o delivers the best open-ended score (49.88%) and the best closed-ended accuracy (58.85%), demonstrating the most effective multimodal reasoning capabilities across task formats.
@@
  • Model scale is not the sole determinant of success. Smaller Qwen (3-7B) and LLaVA (0.5-7B) variants lag behind larger models. However, the Qwen2.5-VL-32B model approximately matches or outperforms its 72B counterpart.
@@
  • Closed-ended questions are easier for MLLMs. Five models achieve an accuracy greater than 50% on closed-ended evaluations (GPT-4o, Claude 3.7 Sonnet, Llama-3.2-Vision-90B, Qwen2.5-VL-32B, Qwen2.5-VL-72B). A text-only baseline still achieves an accuracy of 38.6%, suggesting that multiple-choice items rely less on precise visual grounding than open-ended generation.
  • Expert evaluation validates automated metrics while revealing some differences. Human expert ratings show strong correlation with LLM-as-a-Judge scores in the zero-shot setting (Pearson $r=0.84,p<0.0001$) but only moderate correlation in the ICL setting ($r=0.72,p=0.003$). We find that expert ratings tend to be conservative: medical experts rate some models substantially lower than LLM-as-a-Judge in ICL settings. For example, Qwen2.5-VL-32B and Qwen2.5-VL-72B receive expert ratings of 31-33% despite LLM-as-a-Judge scores of 41-43% in the ICL setting. This suggests that LLM-as-a-Judge may be overly lenient in ICL settings, potentially accepting responses that match the format and phrasing of in-context examples without ensuring clinical adequacy. The high inter-rater agreement among experts ($\geq 98.2\%$) demonstrates that their ratings provide crucial complementary signal to our automated metrics.
Table 1: We benchmark 15 MLLMs on SMMILE, reporting zero-shot performance and performance with in-context examples. The best result is bolded for each task and evaluation metric. Text-only baseline used Llama 3.3 70B. ∗∗Llava-Med-7B refers to LLaVA-Med-v1.5-Mistral-7B. Zero-shot EM scores were $\leq 3.65\%$ for all models and are omitted.